# Condensed Sets and the Solovay Model

### v1

@@ Abstract
  • A Boolean complete topos gives rise to a Boolean-valued model of set theory (potentially with symmetries) by considering its smallest exponential subvariety (its version of the cumulative hierarchy).
@@ 1 Introduction
Theorem (Clausen–Scholze, [14, Sessions 6 and 8]).
@@
The first result was given a forcing theoretic proof in [6] by Bergfalk–Lambie-Hanson–Šaroch, using the interpretation of condensed sets as “generically invariant” information, or more accurately as a system of names over every forcing extension. This insight is based on the classical duality between continuous maps and names, namely that continuous maps from the Stone space of a complete Boolean algebra $\mathbb{B}$ into a compact Hausdorff space $X$ correspond to names for points of the Stone–Čech compactification of $X$ (also called the compact reinterpretation of $X$) in the forcing extension by $\mathbb{B}$:
${\Big\{}\text{continuous }f\colon S(\mathbb{B})\to X{\Big\}}$${\Big\{}\text{names }\tau\in V^{\mathbb{B}},1\Vdash_{\mathbb{B}}\tau\in\beta\check{X}{\Big\}}$}}
Taking this analogy further, Bergfalk–Lambie-Hanson–Šaroch construct a language of propositions that translate between the internal language of the condensed sets and forcing with a specific Boolean algebra $\mathbb{B}$. In an unpublished note, Reid Barton [4] has shown that this translation can be implemented by a geometric morphism into the condensed sets from the category of sheaves on $\mathbb{B}$ with the double negation topology. We may equivalently define this as the category of $\mathbb{B}$-names with names for maps up to forced equality (for a proof see Blass–Scedrov [10, page 44], who attribute it to an unpublished result of Higgs).
@@ 2.1 Stone Duality and Iliadis Absolutes
    for $p\in\mathbb{B}$. This space is Stone, and any homomorphism $h\colon\mathbb{B}\to\mathbb{B}^{\prime}$ gives rise to a continuous map $S(\mathbb{B}^{\prime})\to S(\mathbb{B})$ given again by the preimage $S(h)(U)=h^{-1}U$.
@@ 2.2 The Sites
We refer the reader to [32, Chapter III, Exercise 3] for a list of the coverage axioms. The presence of isomorphisms is clear, as is compositionality for the first site. For the second site, it follows from noting that the forward image of a dense open set under an open map with dense image is dense open.
@@
We note that the above is really a proof of the amalgamation property for complete Boolean algebras, but conducted in the dual category. The dual proof may be found in Monro [34], who shows also a stronger form of the amalgamation property. Many amalgamation properties of Boolean algebras have immediate counterparts in the theory of $Sh(\mathbf{Sol})$, see for instance the proof of Proposition 3.9.
@@
Several remarks are in order about the category $\mathbf{VD}$. First, we should, as usual, be careful about defining definability, but since we always allow parameters from the ground model, the usual tricks with the reflection theorem apply; see [30, Theorem 7.4]. Second, the notation $\mathbf{VD}$ hides the restriction we have made to $V(\mathbb{R})$. Specifically, we are not considering all $V$-definable subsets of $V^{\mathbb{K}}$, but instead those $V$-definable sets that are also hereditarily definable from reals and parameters in $V$. It is also helpful to note the following fact, which was pointed out to us by Asaf Karagila, and simplifies certain aspects of the analysis.
@@
The main fact about this model that we will be using is the following, which can be obtained by minor variations of the argument in Solovay [44]. An exposition of this argument may be found in Jech [25, Theorem 26.14] with Solovay’s original model, or derived from Ikegami [24, Corollary 1.6], whose definition of the Solovay model matches ours.
@@ 2.3 Automorphisms, Names for Generics, and the Like
A standard fact is that if $\mathbb{B}$ is homogeneous, then for all $p\in\mathbb{B}$, and all formulae $\varphi$ with parameters from the ground model, $p\Vdash\varphi$ if and only if $1_{\mathbb{B}}\Vdash\varphi$. The collapse algebra $\mathbb{K}$ satisfies an extremely strong form of homogeneity; see, for instance, [25, Theorem 26.12] for a proof of the following proposition.
@@ 2.4 Geometric Morphism Theory Crash Course
obtained by precomposing with $h$. Since limits and colimits in presheaf categories are taken objectwise, this functor is both complete and cocomplete, and since $\mathbf{A}$ and $\mathbf{B}$ are small, by the adjoint functor theorem $h^{*}$ permits both a left and right adjoint, which turn out to be the left and right Kan extensions respectively. $(h^{*},Ran_{h})$ is thus a geometric morphism between the presheaf topoi where $h^{*}$ has a further left adjoint. This setup arises often in nature, and we will encounter it again in section 3.1. We also set out a catalogue of notions pertaining to geometric morphisms which will be of use.
@@
  • A surjective geometric morphism is one for which $f^{*}$ is faithful. In particular, what will be called a cover of $\mathbf{S}_{1}\to\mathbf{S}_{2}$ is really the inclusion of $\mathbf{S}_{2}$ in $\mathbf{S}_{1}$ as a subcategory.
  • An essential geometric morphism is one for which $f^{*}$ possesses a further left adjoint (denoted $f_{!}$). Note that this is equivalent to $f^{*}$ preserving all set-sized limits.
@@ 3.1 $Sh(\mathbf{Sol})$ covers $\mathbf{Pyk}$ Very Well
There is an essential localic surjection from $Sh(\mathbf{Sol})$ to $\mathbf{Pyk}$ whose inverse image has a further left adjoint which preserves products.
@@
Since $i$ reflects covers, $i$ gives rise to a comorphism of sites [32, Theorem VII.10.5], hence $Ran_{i}$ descends to sheaves (with left adjoint the composite of $i^{*}$ followed by sheafification). We will now show, however, that $i$ is well enough behaved for both $i^{*}$ and $Lan_{i}$ to descend to the sheaf categories.
@@
Since $E$ is the fiber product of Stone spaces, we conclude that the $x_{i}$ form a matching family in $\mathbf{Pyk}$. ∎
@@
  • $Lan_{i}F(\{\})=1$ is immediate ($1$ is still final among complete morphisms).
@@
Since $i^{*}$ has a left and right adjoint, it is complete and cocomplete (respectively), hence in particular $(i^{*},Ran_{i})$ are the left and right part of a geometric morphism $Sh(\mathbf{Sol})\to\mathbf{Pyk}$. We now show that this morphism is quite special, namely that:
@@
  • Every element of $Sh(\mathbf{Sol})$ is a subquotient of one in the image of $i^{*}$ ($(i^{*},Ran_{i})$ is localic).
  • $Lan_{i}$ preserves products.
@@
The proof that $Lan_{i}$ preserves products is more involved, so we decompose it further.
@@
However since $X\to 1$ is open, $f]=[S(\mathbb{B})\to 1]$, hence $Lan_{i}(1)\simeq 1$. ∎
We now turn to showing that the natural map
$Lan_{i}(F\times G)\to Lan_{i}(F)\times Lan_{i}(G)$
is an isomorphism. Note the colimits used to compute both, at the space $S(\mathbb{B})$. We have first
$Lan_{i}(F\times G)(S(\mathbb{B}))=\mathop{\textup{colim}}_{S(\mathbb{B})\to X}(F(X)\times G(X))$
where the colimit is taken over the category of maps
[…]
where $h$ is open and the other two are merely continuous. This is the slice category
$\mathscr{S}=const_{S(\mathbb{B})}\downarrow i.$
In turn, the right hand side is given by
$\mathop{\textup{colim}}_{S(\mathbb{B})\to X}\mathop{\textup{colim}}_{S(\mathbb{B})\to Y}(F(X)\times G(Y))$
which is the colimit over the product category $\mathscr{S}\times\mathscr{S}$. Since $F$ and $G$ are contravariant, we wish to show that the diagonal functor $\mathscr{S}^{op}\to\mathscr{S}^{op}\times\mathscr{S}^{op}$ is final. To show this, we note the result of Adámek–Rosický [1](cf. also Gabriel–Ulmer [19]), showing that this is equivalent to $\mathscr{S}^{op}$ being sifted.
Definition 3.8.
A category is sifted if for every pair of its objects $c,d$, the category of cospans $cospan(c,d)$ is connected.
The opposite category $\mathscr{S}^{op}$ is more appropriately a category of Boolean algebras, so we shift to algebraic thinking, and set out some notation. For the purpose of this proof,
  • Greek letters denote complete homomorphisms,
  • Latin letters denote not-necessarily-complete homomorphisms,
  • $\overline{(\cdot)}$ denotes, as before, the Dedekind–Macneille completion,
  • $\cdot\otimes\cdot$ denotes, as in definition 2.9, the tensor product.
Now the siftedness of $\mathscr{S}^{op}$ follows from the following (note that the statement is a priori stronger than siftedness).
@@
Fix the complete Boolean algebra $\mathbb{B}$, and a pair of (potentially non-complete) morphisms $c\colon\mathbb{C}\to\mathbb{B},d\colon\mathbb{D}\to\mathbb{B}\in\mathscr{S}^{op}$ for complete algebras $\mathbb{C}$ and $\mathbb{D}$. Then we have that
  • There exists some complete Boolean algebra $U$, complete morphisms $\mu_{\mathbb{C}}$ and $\mu_{\mathbb{D}}$, and a non-complete morphism $u$ completing the commutative diagram below (note the directions)
    ${U}{\mathbb{C}}$$\mathbb{D}$$\mathbb{B}$$\scriptstyle{\mu_{\mathbb{C}}}$$\scriptstyle{c}$$\scriptstyle{\mu_{\mathbb{D}}}$$\scriptstyle{d}$$\scriptstyle{u}$
    (We call such a diagram an amalgam of $\mathbb{C}$ and $\mathbb{D}$ with markings $c$ and $d$.)
  • For any amalgam $U$, there is a complete collapse algebra $K=\overline{Coll(\omega,\alpha)}$ with $2^{|\mathbb{B}|}<\alpha<\kappa$ so that there is a “map of amalgams” from $U$ into $K$, i.e. there are complete morphisms $\nu$, $\mu^{\prime}_{\mathbb{C}}$, $\mu^{\prime}_{\mathbb{D}}$ and a morphism $k$ which fit into the commutative diagram
[…]
  • For any two complete algebras $K_{1}=\overline{Coll(\omega,\alpha_{1})}$, $K_{2}=\overline{Coll(\omega,\alpha_{2})}$ with $2^{|\mathbb{B}|}<\alpha_{1},\alpha_{2}<\kappa$ that are amalgams of $\mathbb{C}$ and $\mathbb{D}$ with markings $c$ and $d$,
    $K=\overline{K_{1}\times_{\mathbb{B}}K_{2}}$
    is a “common refinement”, i.e. there are complete morphisms $\mu_{\mathbb{C}}$, $\mu_{\mathbb{D}}$, $\nu_{1}$, $\nu_{2}$, and a morphism $k$ which fit into the commutative diagram
[…]
In particular, the cospan category $cospan(c,d)$ is connected, hence $\mathscr{S}^{op}$ is sifted.
Proof.
We prove the parts in order, highlighting the observations that do not follow from categorical generalities in bold.
  • Set $U=\overline{\mathbb{C}\otimes\mathbb{D}}$ with $\mu_{\mathbb{C}}$, $\mu_{\mathbb{D}}$ the embedding maps. Crucially, these are complete (dually, the projection maps out of a product are open). Define $u$ in the natural way, i.e.
    $u(a\otimes b)=c(a)\wedge d(b)$
    which is a not necessarily complete homomorphism $\mathbb{C}\otimes\mathbb{D}\to\mathbb{B}$, hence lifts to $U$ by the completeness of $\mathbb{B}$.
  • Choose any $\alpha<\kappa$ sufficiently large so that $U$ embeds as a complete subalgebra of $K=\overline{Coll(\omega,\alpha)}$ (using, for instance, that $\kappa$ is strong limit). Let $\nu\colon U\to K$ be this embedding, and define $\mu^{\prime}_{\mathbb{C}}=\nu\circ\mu_{\mathbb{C}}$, $\mu^{\prime}_{\mathbb{D}}=\nu\circ\mu_{\mathbb{D}}$, which are complete due to being composites of complete maps. To obtain $k$, note the not-necessarily-complete morphism
    $u\circ\nu^{-1}\colon\nu[U]\to\mathbb{B}.$
    This corresponds to a $\mathbb{B}$-name for a not-necessarily-complete ultrafilter $F$ on $\nu[U]\subseteq K$, which is thus a filter on $K$. In the forcing extension by $\mathbb{B}$, use the ultrafilter principle to take an ultrafilter $F^{\prime}$ of $K$ extending $F$. The name for $F^{\prime}$ over $\mathbb{B}$ (which exists by the maximality principle) then corresponds to a map $K\to\mathbb{B}$ that extends $u\circ\nu^{-1}$. This is $k$, which manifestly satisfies $k\circ\nu=u$.
  • First we show the existence of the dotted homomorphisms, and then (the hard part) verify that they are complete. The $\nu_{i}\colon K\to K_{i}$ are the projection maps of the pullback, and $k=k_{1}\circ\nu_{1}=k_{2}\circ\nu_{2}$ (which are equal by definition of the pullback). Moreover, since
    $k_{1}\circ\mu^{1}_{\mathbb{C}}=k_{2}\circ\mu^{2}_{\mathbb{C}}=c,$
    there exists a map $\mu_{\mathbb{C}}\colon\mathbb{C}\to K$ by the universal property of the pullback, which further satisfies $\nu_{i}\circ\mu_{\mathbb{C}}=\mu^{i}_{\mathbb{C}}$. Similarly, since
    $k_{1}\circ\mu^{1}_{\mathbb{D}}=k_{2}\circ\mu^{2}_{\mathbb{D}}=d$
    we have a map $\mu_{\mathbb{D}}\colon\mathbb{D}\to K$ satisfying $\nu_{i}\circ\mu_{\mathbb{D}}=\mu^{i}_{\mathbb{D}}$. It remains to check that $\nu_{1},\nu_{2}$ and $\mu_{\mathbb{C}}$, $\mu_{\mathbb{D}}$ are complete. Note that if the $\nu_{i}$ are complete, then the left and right projections of $\mu_{\mathbb{C}}$ and $\mu_{\mathbb{D}}$ ($\mu^{i}_{\mathbb{C}}$ and $\mu^{i}_{\mathbb{D}}$ respectively) are complete, hence so are $\mu_{\mathbb{C}}$ and $\mu_{\mathbb{D}}$. Thus it remains to check the completeness of $\nu_{1}$ (the proof for $\nu_{2}$ being the same). We will show completeness at the level of $K_{1}\times_{\mathbb{B}}K_{2}$, as this suffices (complete maps extend to complete maps, for instance see Givant and Halmos [20, Chapter 25, Exercise 12]).
    The crucial point is that the kernel of $k_{1}$ is dense, since every $K_{1r}$ has the same cardinality (by ultrahomogeneity), and so $k_{1}$ restricted to $K_{1r}\to\mathbb{B}_{k_{1}(r)}$ cannot be injective, since $|\mathbb{B}|<|K_{1}|$. Specifically, every non-unit $r$ has $1>r^{\prime}\geq r$ so that $k_{1}(r^{\prime})=1$.
    Now, let $(p_{\alpha},q_{\alpha})$ be a family of pairs with supremum $(1,1)\in K_{1}\times_{\mathbb{B}}K_{2}$, i.e.
    $(p,q)\geq\left(\bigvee_{\alpha}p_{\alpha},\bigvee_{\alpha}q_{\alpha}\right)\wedge k_{1}(p)=k_{2}(q)\rightarrow p=1,q=1$
    If $\bigvee_{\alpha}p_{\alpha}$ were not $1$, there would be $1>p\geq\bigvee_{\alpha}p_{\alpha}$ with $k_{1}(p)=1$, but then the pair $(p,1)$ is a contradiction to the previous sentence. Hence $\bigvee_{\alpha}p_{\alpha}=1$, but this shows that the left projection is complete. Hence $\nu_{1}$, its lift, is complete.
Thus we have completed the proof that $Lan_{i}$ preserves products, and hence of our theorem. We note for completeness that $Lan_{i}$ does not preserve equalisers, hence does not preserve pullbacks either.
Proposition 3.10.
@@
Let $X$ be extremally disconnected compact Hausdorff that lacks isolated points, and let $Y=\mathscr{A}(X\times X)$ be the absolute of its products, with its two open projection maps

### v2

@@ Abstract
  • A Boolean complete topos gives rise to a symmetric Boolean-valued model of set theory by considering its smallest exponential subvariety (its version of the cumulative hierarchy).
@@ 1 Introduction
Theorem (Clausen–Scholze, [13, Sessions 6 and 8]).
@@
The first result was given a forcing theoretic proof in [5] by Bergfalk–Lambie-Hanson–Šaroch, using the interpretation of condensed sets as “generically invariant” information, or more accurately as a system of names over every forcing extension. This insight is based on the classical duality between continuous maps and names, namely that continuous maps from the Stone space of a complete Boolean algebra $\mathbb{B}$ into a compact Hausdorff space $X$ correspond to names for points of the compact reinterpretation of $X$ in the forcing extension by $\mathbb{B}$:
${\Big\{}\text{continuous }f\colon S(\mathbb{B})\to X{\Big\}}$${\Big\{}\text{names }\tau\in V^{\mathbb{B}},1\Vdash_{\mathbb{B}}\tau\in\overline{X}{\Big\}}$},}
where $\overline{X}$ denotes the set of $V^{\mathbb{B}}$ maximal filters of closed $V$-subsets of $X$. Taking this analogy further, Bergfalk–Lambie-Hanson–Šaroch construct a language of propositions that translate between the internal language of the condensed sets and forcing with a specific Boolean algebra $\mathbb{B}$. In an unpublished note, Reid Barton [3] has shown that this translation can be implemented by a geometric morphism into the condensed sets from the category of sheaves on $\mathbb{B}$ with the double negation topology. We may equivalently define this as the category of $\mathbb{B}$-names with names for maps up to forced equality – for a proof see Blass–Scedrov [8, page 44], who attribute it to an unpublished result of Higgs.
@@ 2.1 Stone Duality and Iliadis Absolutes
    for $p\in\mathbb{B}$. This space is Stone, and any homomorphism $h\colon\mathbb{B}\to\mathbb{C}$ gives rise to a continuous map $S(\mathbb{C})\to S(\mathbb{B})$ given again by the preimage $S(h)(U)=h^{-1}U$.
@@ 2.2 The Sites
We refer the reader to [29, Chapter III, Exercise 3] for a list of the coverage axioms. The presence of isomorphisms is clear, as is compositionality for the first site. For the second site, it follows from noting that the forward image of a dense open set under an open map with dense image is dense open.
@@
We note that the above is really a proof of the amalgamation property for complete Boolean algebras, but conducted in the dual category. The dual proof may be found in Monro [32], who shows also a stronger form of the amalgamation property. Many amalgamation properties of Boolean algebras seem to have immediate counterparts in the theory of $Sh(\mathbf{Sol})$.
@@
Several remarks are in order about the category $\mathbf{VD}$. First, we should, as usual, be careful about defining definability, but since we always allow parameters from the ground model, the usual tricks with the reflection theorem apply; see [28, Theorem 7.4]. Second, the notation $\mathbf{VD}$ hides the restriction we have made to $V(\mathbb{R})$. Specifically, we are not considering all $V$-definable subsets of $V^{\mathbb{K}}$, but instead those $V$-definable sets that are also hereditarily definable from reals and parameters in $V$. It is also helpful to note the following fact, which was pointed out to us by Asaf Karagila, and simplifies certain aspects of the analysis.
@@
The main fact about this model that we will be using is the following, which can be obtained by minor variations of the argument in Solovay [42]. An exposition of this argument may be found in Jech [23, Theorem 26.14] with Solovay’s original model, or derived from Ikegami [22, Corollary 1.6], whose definition of the Solovay model matches ours.
@@ 2.3 Automorphisms, Names for Generics, and the Like
A standard fact is that if $\mathbb{B}$ is homogeneous, then for all $p\in\mathbb{B}$, and all formulae $\varphi$ with parameters from the ground model, $p\Vdash\varphi$ if and only if $1_{\mathbb{B}}\Vdash\varphi$. The collapse algebra $\mathbb{K}$ satisfies an extremely strong form of homogeneity; see, for instance, [23, Theorem 26.12] for a proof of the following proposition.
@@ 2.4 Geometric Morphism Theory Crash Course
obtained by precomposing with $h$. Since limits and colimits in presheaf categories are taken objectwise, this functor is both continuous and cocontinuous, and since $\mathbf{A}$ and $\mathbf{B}$ are small, by the adjoint functor theorem $h^{*}$ permits both a left and right adjoint, which turn out to be the left and right Kan extensions respectively. $(h^{*},Ran_{h})$ is thus a geometric morphism between the presheaf topoi where $h^{*}$ has a further left adjoint. This setup arises often in nature, and we will encounter it again in section 3.1. We also set out a catalogue of notions pertaining to geometric morphisms which will be of use.
@@
  • A surjective geometric morphism is one for which $f^{*}$ is faithful. In particular, what will be called a cover $\mathbf{S}_{1}\to\mathbf{S}_{2}$ is really the inclusion of $\mathbf{S}_{2}$ in $\mathbf{S}_{1}$ as a subcategory.
  • A terminally connected geometric morphism is one for which $f^{*}$ possesses a further left adjoint $f_{!}$ which furthermore preserves the terminal object ($f_{!}(1)\simeq 1$).
@@ 3.1 $Sh(\mathbf{Sol})$ covers $\mathbf{Pyk}$ Very Well
There is a terminally connected localic surjection from $Sh(\mathbf{Sol})$ to $\mathbf{Pyk}$.
@@
Since $i$ reflects covers, $i$ gives rise to a comorphism of sites [29, Theorem VII.10.5], hence $Ran_{i}$ descends to sheaves (with left adjoint the composite of $i^{*}$ followed by sheafification). We will now show, however, that $i$ is well enough behaved for both $i^{*}$ and $Lan_{i}$ to descend to the sheaf categories.
@@
Since $E$ is the fiber product of Stone spaces, we conclude that the $x_{i}$ form a matching family for the site $\mathbf{Stone}$. In particular, $F$ has a unique gluing for this family. ∎
@@
  • $Lan_{i}F(\{\})=1$ is immediate (the empty map is open).
@@
Since $i^{*}$ has a left and right adjoint, it is continuous and cocontinuous (respectively), hence in particular $(i^{*},Ran_{i})$ are the left and right part of a geometric morphism $Sh(\mathbf{Sol})\to\mathbf{Pyk}$. We now show that this morphism is quite special, namely that:
@@
  • Every object of $Sh(\mathbf{Sol})$ is a subquotient of one in the image of $i^{*}$ ($(i^{*},Ran_{i})$ is localic).
  • $Lan_{i}$ preserves the terminal object ($(i^{*},Ran_{i})$ is terminally connected).
@@
However since $X\to 1$ is open, $S(\mathbb{B})\xrightarrow{f}X]=[S(\mathbb{B})\xrightarrow{f}X\to 1]$, hence $Lan_{i}(1)\simeq 1$. ∎
This completes the proof of Theorem 3.1. It is reasonable to ask whether more can be shown about this morphism. We note for completeness that $Lan_{i}$ does not preserve other limits, such as equalisers. [^1] An earlier draft of this paper asserted that $Lan_{i}$ preserved products, but the proof contained an error. See Lemma B.3 for a more modest claim that can be shown.
@@
Let $X$ be extremally disconnected compact Hausdorff that lacks isolated points, and let $Y=\mathscr{A}(X\times X)$ be the absolute of its product with itself, with its two open projection maps